Business Analytics Harrys Hardware does a brisk husivess during the yeasr, bun daringCsea sells Christmas trees for a subetanial profit. Unfonu
Answer #1

Demand PDemand - ら。 0.05 :00 75 0.aS (00 0-20 0.85 25 0 30 .20 O 35 0 15 오。 0.05 0.05 Hし:20 HP=80-20-60 = 0, 25 20 60 Stock 5CCO) Demand PlDemahd) 50 0-25 I+ 00 75 0 25 0.25 O.25 0-75 (oO 0. 50 125 0.25 HL=20; NP=100-20 HLI HL MP) (2080 O.20 Stock 12